Circular No. 6742

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dd IN NGC 6060 P. Garnavich and S. Jha, Harvard-Smithsonian Center for Astrophy-

sics; and R. Kirshner, University of California at Santa Barbara, report: “Spectra of SN 1997dd (cf. IAUC 6724) were obtained by P. Berlind with the 1.5-m Tillinghast telescope on Sept. 10.1 UT. Strong absorption lines have developed at 581.2, 660.0, and 699.1 nm, corresponding to He I with a photospheric velocity of 8000 km/s. The photospheric velocity of Hα is 12 000 km/s. The spectrum is similar to that of SN 1996cb ∼ 1 month after outburst and confirms the prediction by Phillips and Suntzeff (IAUC 6725) that SN 1997dd would transform into a type-Ib supernova.”

GRB 970815 J. Greiner, Astrophysical Institute, Potsdam, reports: “A 17 000-s

ROSAT observation of the RXTE error box of GRB 970815 (cf. IAUC 6718) during Aug. 20–22 reveals ten x-ray sources within the HRI field-of-view. One faint source is detected at α = 16^{h}08^{m}47^{s}.7, δ = +81^{o}31′53′′ (equinox 2000.0), which just falls on the border of the RXTE error box. Using a power-law model with photon index 2 and the galactic absorbing column N_{H} = 4.7 × 10^{20} cm^{−2}, the observed countrate of 0.00073 ± 0.00025 counts/s corresponds to an unabsorbed ROSAT flux (band 0.1– 2.4 keV) of 5 × 10^{−14} erg cm^{−2} s^{−1}. This source is too faint to make any variability analysis. However, its flux level is consistent with a t^{−1} flux decay, suggesting that it may be the afterglow of GRB 970815. One of the x-ray sources detected outside the RXTE error box is located at α = 16^{h}06^{m}51^{s}.5, δ = +81^{o}30′28′′, consistent with the ASCA detection (IAUC 6722). Its countrate of 0.0035 ± 0.0005 counts/s corresponds to an extrapolated flux in the band 2–10 keV that is comparable within a factor of 2 with the ASCA source flux. No optical object is found within the 10′′ error boxes of both x-ray sources on the digitized POSS and on Whipple Observatory images (IAUC 6721). The x-ray image is posted at

http://www.aip.de:8080/~jcg/grb970815.html.”

COMET C/1997 N1 (TABUR) S. Nakano, Sumoto, Japan, reports that H. Abe (Yatsuka) was unable

to locate this comet (cf. IAUC 6714) on several CCD images obtained with a 0.26-m reflector around Sept. 8.44 UT, though stars of mag 15 were visible near the comet’s predicted position.

1997 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green